Simplify and solve the following linear equations. 
 15(y – 4) –2(y – 9) + 5(y + 6) = 0 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

15(y–4)–2(y–9)+5(y+6)=0
⇒15y-60-2y+18+5y+30=0
⇒15y-2y+5y=60-18-30
⇒18y=12
⇒y=12/18
⇒y=2/3

Thanks